# ApexNav: An Adaptive Exploration Strategy for Zero-Shot Object Navigation with Target-centric Semantic Fusion

### v2

@@ Abstract
Navigating unknown environments to find a target object is a significant challenge. While semantic information is crucial for navigation, relying solely on it for decision-making may not always be efficient, especially in environments with weak semantic cues. Additionally, many methods are susceptible to misdetections, especially in environments with visually similar objects. To address these limitations, we propose ApexNav, a zero-shot object navigation framework that is both more efficient and reliable. For efficiency, ApexNav adaptively utilizes semantic information by analyzing its distribution in the environment, guiding exploration through semantic reasoning when cues are strong, and switching to geometry-based exploration when they are weak. For reliability, we propose a target-centric semantic fusion method that preserves long-term memory of the target object and similar objects, reducing false detections and minimizing task failures. We evaluate ApexNav on the HM3Dv1, HM3Dv2, and MP3D datasets, where it outperforms state-of-the-art methods in both SR and SPL metrics. Comprehensive ablation studies further demonstrate the effectiveness of each module. Furthermore, real-world experiments validate the practicality of ApexNav in physical environments. Project page is available at ApexNav.
@@ I Introduction
Nonetheless, current ZSON methods still face challenges in efficiency and robustness. First, most methods [24, 25, 26, 28] rely heavily on semantic reasoning to explore and search. However, semantic guidance is not always available or reliable. For example, at the beginning of task, the agent may face a blank wall or occlusions, limiting semantic cues and hindering effective semantic-based navigation. Moreover, ambiguous or weak semantic cues can mislead the agent. For example, a plant might appear in various rooms like a bedroom, living room, or bathroom, and on different surfaces like a table, windowsill, or floor, making it hard to infer its location without sufficient context. In summary, over-reliance on semantics can reduce navigation efficiency. Secondly, some approaches [6, 25] rely on single-frame detections to identify targets, while others [8, 26] adopt max-confidence fusion to build semantic maps for target identification. However, relying solely on single-frame results often lacks robustness. The max-confidence fusion approach also struggles with misdetections, as even sporadic high-confidence misdetections can result in incorrect semantic labels that are hard to correct through subsequent observations. These issues are exacerbated in cluttered scenes with occlusions or similar-looking objects. To improve robustness, multi-frame integration and more reliable fusion strategies are needed.
@@
2) Target-centric semantic fusion: ApexNav maintains long-term memory of the target object and visually similar objects. By aggregating multi-frame observations with context-aware confidence weighting, it improves robustness to misclassification and enhances the accuracy of target identification.
@@ II-B Exploration Strategy for ZSON
Current ZSON methods mainly use frontier-based exploration. CoW [6] follows a nearest-frontier strategy based only on geometry, ignoring semantic cues in indoor environments. L3MVN [26] and TriHelper [28] construct a semantic map and use LLMs to identify target-related frontiers, while SG-Nav [24] leverages 3D scene graphs to prompt LLMs for reasoning. VLFM [25] applies a VLM to build a value map linking frontiers to the target. These methods typically score frontiers with LLMs or VLMs and use a greedy policy to select the highest-scoring one. While effective in semantically rich environments, they often perform poorly when semantic cues are weak. SemUtil [5] notes that early semantic guidance is inefficient and adopts a fixed two-stage strategy that starts with geometry for 50 steps and then switches to semantics. However, this fixed scheme lacks adaptability. To address this, we propose an adaptive strategy that effectively leverages semantic and geometry information to guide exploration based on the environment’s semantic distribution.
@@ II-C Target Identification for ZSON
The success rate of ZSON is directly influenced by accurate target identification, typically using pre-trained detection and segmentation models [11, 19, 27] on RGB inputs. Some methods [6, 25] rely solely on single-frame detector outputs, making them vulnerable to failure from mispredictions. Several works [8, 26] construct semantic maps via max-confidence fusion, which is prone to high-confidence misdetections that are difficult to correct later. To mitigate this, TriHelper [28] and SG-Nav [24] introduce validation modules: the former leverages VLMs for refinement, while the latter combines multi-frame scores with 3D scene graphs. Although these methods improve reliability, they lack a robust fusion mechanism to effectively manage detector errors in complex environments. We propose a target-centric semantic fusion method that aggregates multi-frame observations using context-aware confidence updates, effectively filtering out occasional errors and improving robustness.
@@ IV-B2 Semantic-based Exploration
$\mathbf{M_{tsp}}(i,j)=\begin{cases}0&\text{if }j=0,\\ D(i,j)&\text{if }i\in[0,N_{h}],j\in(0,N_{h}]\end{cases}$ (1)
@@ IV-C1 Object Reasoning with LLMs
We leverage the reasoning ability of LLMs with Chain-of-Thought prompting [21] to identify objects that are easily misclassified as the target. Considering that detector performance can be affected by object size—small objects often yield low confidence—the LLM is also prompted to generate an adaptive confidence threshold $c_{th}$ for the target. An object is treated as a reliable target only if its fused confidence exceeds this threshold. The agent then navigates to the object. This process yields a comprehensive object list $L_{obj}$ with $M_{obj}$ entries, including the target and similar objects, guiding subsequent decisions.
@@ IV-C3 Semantic Label Confidence Fusion
$\begin{split}n_{new}^{l_{det}^{i}}&=n_{last}^{l_{det}^{i}}+v_{det}^{i},\\ c_{new}^{l_{det}^{i}}&=\frac{n_{last}^{l_{det}^{i}}}{n_{new}^{l_{det}^{i}}}\cdot c_{last}^{l_{det}^{i}}+\frac{v_{det}}{n_{new}^{l_{det}^{i}}}\cdot c_{det}^{i},\end{split}$ (2)
@@ IV-D1 Efficiency Cost
$cost_{target}=\omega_{t}\cdot\|p_{new}-p_{local}\|$ (3)
@@
$\begin{split}cost_{prox}=&\omega_{p}\cdot\left(\|p_{new}-p_{local}\|-\|p_{cur}-p_{local}\|\right)\\ &\cdot\mathds{1}_{\|p_{new}-p_{local}\|>\|p_{cur}-p_{local}\|}\end{split}$ (4)
@@
Method Zero-shot HM3Dv1 HM3Dv2 MP3D
[…]
SemExp[4] No 36.0 14.4
PONI[17] No 31.8 12.1
ZSON[13] No 25.5 12.6 15.3 4.8
CoW[6] Yes 7.4 3.7
ESC[30] Yes 39.2 22.3 28.7 14.2
L3MVN[26] Yes 50.4 23.1 36.3 15.7 34.9 14.5
OpenFMNav[8] Yes 54.9 24.4 37.2 15.7
InstructNav[12] Yes 58.0 20.9
VLFM[25] Yes 52.5 30.4 63.6 32.5 36.4 17.5
TriHelper[28] Yes 56.5 25.3
SG-Nav[24] Yes 54.0 24.9 49.6 25.5 40.2 16.0
ApexNav(Ours) Yes 59.6 33.0 76.2 38.0 39.2 17.8
@@ IV-D2 Safety Cost
$\begin{split}cost_{safe}=\sum_{i=1}^{K}\frac{\omega_{s}}{d_{obs}(p_{sa_{i}})+\epsilon}\cdot\mathds{1}_{d_{obs}(p_{sa_{i}})<d_{th}}\end{split}$ (5)
@@ V Experiments
Figure 7: Failure Cause Statistics of ApexNav on the Three Datasets.
@@ V-A Simulation Experimental Setup
Baselines: SemExp [4], PONI [17], and ZSON [13] rely on task-specific training, limiting their generalization in zero-shot settings. CoW [6] uses a nearest-frontier exploration strategy without semantic reasoning. ESC [30], L3MVN [26], TriHelper [28], and OpenFMNav [8] improve exploration by building semantic maps and using LLMs to select frontiers. VLFM [25] and InstructNav [12] use LLMs or VLMs to create value maps that guide agents to targets. SG-Nav [24] further enhance semantic reasoning by prompting LLMs with 3D scene graphs to select frontiers.
@@ V-B Comparison with State-of-the-art
V-C Failure Cause Analysis
Task outcomes are categorized as follows:
Except for the HM3Dv2 dataset, which mainly consists of single-floor tasks, the others show many ”Different Floor” failures ($>13\%$). This is because ApexNav is designed for single-floor task and lacks strategies for cross-floor mapping. This limitation also leads to many ”No Frontier” failures. When the agent starts at a staircase between two floors, the vertical distance to the target is not large enough to be recognized as on a different floor. However, since the agent builds a 2D map centered at its starting position, it may find no frontiers to explore after initialization, resulting in task failure. These problems arise from the fact that a single 2D map cannot fully support exploration in 3D environments.
@@ V-C Failure Cause Analysis
V-D Ablation Study on Adaptive Exploration Strategy
V-D1 Ablation Study on Exploration Strategy
We conducted an ablation study on HM3Dv2 to evaluate how different exploration strategies affect performance. Four strategies were tested: (a) TSP-Semantic: selects high-scoring frontiers and solves TSP to plan the order; (b) Greedy-Semantic: chooses the frontier with the highest semantic score; (c) TSP-Geometry: solves TSP over all frontiers; (d) Greedy-Geometry: selects the nearest frontier. We evaluated both single-stage strategies (using one method throughout) and two-stage strategies (switching between semantic and geometric modes based on thresholds $\sigma_{t}=0.015$ and $r_{t}=1.10$). Results are reported in SR, SPL, and SoftSPL, as shown in Table II.
@@ V-D1 Ablation Study on Exploration Strategy
The results show that two-stage strategies outperform single-stage ones, highlighting the advantage of adaptively switching between semantic and geometric exploration. Among the single-stage strategies, TSP-Semantic and Greedy-Geometry performed best, and their combination achieves the highest overall score: Greedy-Geometry is efficient when semantic cues are weak, while TSP-Semantic guides the agent effectively once strong semantic cues emerge.
@@ V-D2 Ablation Study on Exploration Metrics
V-E Ablation Study on Target-centric Semantic Fusion
To assess the impact of ApexNav’s target-centric semantic fusion method on SR and SPL, we conducte ablation experiments on HM3Dv1, HM3Dv2, and MP3D. The following control groups were designed: (a) ApexNav w/o Similar Objects: Removes the auxiliary detection module for similar objects, focusing only on detecting the target; (b) ApexNav w/o Fusion: Removes the semantic fusion module, directly overwriting confidence scores with detection results; (c) ApexNav w/o Observation: The confidence of the object cluster is not reduced by fusing low-confidence detection results when no objects are detected; (d) ApexNav w Max-fusion: Replaces ApexNav’s fusion method with max-confidence fusion. The experimental results are presented in Table IV.
The (b) group without fusion shows the worst performance, highlighting the importance of fusion. Additionally, ApexNav outperforms the max-fusion approach (d), proving its effectiveness in preventing task failures caused by high-confidence mislabeling. The other two experiments emphasize the significance of fusing similar objects and reducing confidence for objects that were previously detected but are currently undetected. Overall, the target-centric semantic fusion method delivers the best performance, demonstrating its effectiveness.
@@ V-E Ablation Study on Target-centric Semantic Fusion
w/o Similar Objects 56.95 30.40 75.00 37.63 28.70 15.43
w/o Fusion 44.50 25.48 55.00 29.72 27.84 14.03
w/o Observation 50.48 29.17 70.20 37.44 35.58 16.09
w Max-fusion 51.43 28.97 67.90 37.04 32.11 15.32
@@
V-F Real-world Deployment
We conduct real-world experiments using the AgileX LIMO robot. A laptop equipped with an RTX 4060 Laptop GPU and an AMD R9 7945HX CPU served as the ROS master and communicated with the robot via a ROS-based master-slave architecture. We use MINCO [20] to generate collision-free spatiotemporal trajectories for waypoint navigation. Semantic segmentation and BLIP-2 run at 3 Hz. We perform extensive experiments across various object goals in both home and laboratory environments. All trials successfully locate the target objects, demonstrating the feasibility of ApexNav in real-world settings. Representative results are shown in Fig. 1.

### v3

@@ Abstract
Navigating unknown environments to find a target object is a significant challenge. While semantic information is crucial for navigation, relying solely on it for decision-making may not always be efficient, especially in environments with weak semantic cues. Additionally, many methods are susceptible to misdetections, especially in environments with visually similar objects. To address these limitations, we propose ApexNav, a zero-shot object navigation framework that is both more efficient and reliable. For efficiency, ApexNav adaptively utilizes semantic information by analyzing its distribution in the environment, guiding exploration through semantic reasoning when cues are strong, and switching to geometry-based exploration when they are weak. For reliability, we propose a target-centric semantic fusion method that preserves long-term memory of the target and similar objects, enabling robust object identification even under noisy detections. We evaluate ApexNav on the HM3Dv1, HM3Dv2, and MP3D datasets, where it outperforms state-of-the-art methods in both SR and SPL metrics. Comprehensive ablation studies further demonstrate the effectiveness of each module. Furthermore, real-world experiments validate the practicality of ApexNav in physical environments. The code will be released at https://github.com/Robotics-STAR-Lab/ApexNav.
@@ I Introduction
Nonetheless, current ZSON methods still face challenges in efficiency and robustness. First, most methods [26, 27, 28, 30] rely heavily on semantic reasoning to explore and search. However, semantic guidance is not always available or reliable. For example, at the beginning of task, the agent may face a blank wall or occlusions, limiting semantic cues and hindering effective semantic-based navigation. Moreover, ambiguous or weak semantic cues can mislead the agent. For example, a plant might appear in various rooms like a bedroom, living room, or bathroom, and on different surfaces like a table, windowsill, or floor, making it hard to infer its location without sufficient context. In summary, over-reliance on semantics can reduce navigation efficiency. Secondly, some approaches [7, 27] rely on single-frame detections to identify targets, while others [9, 28] adopt max-confidence fusion to build semantic maps for target identification. However, relying solely on single-frame results often lacks robustness. The max-confidence fusion approach also struggles with misdetections, as even sporadic high-confidence misdetections can result in incorrect semantic labels that are hard to correct through subsequent observations. These issues are exacerbated in cluttered scenes with occlusions or similar-looking objects. Therefore, improving robustness under noisy detection requires multi-frame integration and more reliable fusion strategies.
@@
2) Target-centric semantic fusion: ApexNav maintains long-term memory of the target object and visually similar objects. By aggregating multi-frame observations with context-aware confidence weighting, it enables reliable target identification even when using noisy and imperfect detectors.
@@ II-B Exploration Strategy for ZSON
Current ZSON methods mainly use frontier-based exploration. CoW [7] follows a nearest-frontier strategy based only on geometry, ignoring semantic cues in indoor environments. L3MVN [28] and TriHelper [30] construct a semantic map and use LLMs to identify target-related frontiers, while SG-Nav [26] leverages 3D scene graphs to prompt LLMs for reasoning. VLFM [27] applies a VLM to build a value map linking frontiers to the target. These methods typically score frontiers with LLMs or VLMs and use a greedy policy to select the highest-scoring one. WTRP-Searcher [12] instead models ObjectNav as a Weighted Traveling Repairman Problem to overcome greedy policy limitations. While effective in semantically rich environments, they often perform poorly when semantic cues are weak. SemUtil [6] notes that early semantic guidance is inefficient and adopts a fixed two-stage strategy that starts with geometry for 50 steps and then switches to semantics. However, this fixed scheme lacks adaptability. Moreover, SemUtil ranks frontiers using a utility function (semantic score divided by path length), the trade-off between semantic relevance and distance may suppress promising frontiers and limit overall performance. To address these, we propose an adaptive strategy that effectively leverages semantic and geometry information to guide exploration based on the environment’s semantic distribution.
@@ II-C Target Identification for ZSON
The success rate of ZSON is directly influenced by accurate target identification, typically using pre-trained detection and segmentation models [13, 21, 29] on RGB inputs. As noted in [3], false positives from these detectors remain a major limitation. Some methods [7, 27] rely solely on single-frame detector outputs, making them vulnerable to failure from mispredictions. Several works [9, 28] construct semantic maps via max-confidence fusion, which is prone to high-confidence misdetections that are difficult to correct later. To mitigate this, TriHelper [30] and SG-Nav [26] introduce validation modules: the former leverages VLMs for refinement, while the latter combines multi-frame scores with 3D scene graphs. Although these methods improve reliability, they lack a robust fusion mechanism to effectively manage detector errors in complex environments. We propose a target-centric semantic fusion method that aggregates multi-frame observations using context-aware confidence updates, effectively filtering out occasional errors and improving robustness.
@@ IV-B2 Semantic-based Exploration
[…]
@@ IV-C1 Object Reasoning with LLMs
We leverage the reasoning ability of LLMs with Chain-of-Thought prompting [23] to identify objects that are easily misclassified as the target. Considering that detector performance can be affected by object size, as small objects often yield low confidence, the LLM is also prompted to generate an adaptive confidence threshold $c_{th}$ for the target. An object is treated as a reliable target only if its fused confidence exceeds this threshold. The agent then navigates to the object. This process yields a comprehensive object list $L_{obj}$ with $M_{obj}$ entries, including the target and similar objects, guiding subsequent decisions.
@@ IV-C3 Semantic Label Confidence Fusion
$\begin{split}n_{new}^{l_{det}^{i}}&=n_{last}^{l_{det}^{i}}+v_{det}^{i},\\ c_{new}^{l_{det}^{i}}&=\frac{n_{last}^{l_{det}^{i}}}{n_{new}^{l_{det}^{i}}}\cdot c_{last}^{l_{det}^{i}}+\frac{v_{det}}{n_{new}^{l_{det}^{i}}}\cdot c_{det}^{i}.\end{split}$ (2)
@@ IV-D1 Efficiency Cost
$cost_{target}=\omega_{t}\cdot\|p_{new}-p_{local}\|.$ (3)
@@
$\begin{split}cost_{prox}=&\omega_{p}\cdot\left(\|p_{new}-p_{local}\|-\|p_{cur}-p_{local}\|\right).\\ &\cdot\mathds{1}_{\|p_{new}-p_{local}\|>\|p_{cur}-p_{local}\|}\end{split}$ (4)
@@ IV-D2 Safety Cost
$\begin{split}cost_{safe}=\sum_{i=1}^{K}\frac{\omega_{s}}{d_{obs}(p_{sa_{i}})+\epsilon}\cdot\mathds{1}_{d_{obs}(p_{sa_{i}})<d_{th}}.\end{split}$ (5)
@@ V-A Simulation Experimental Setup
SemExp[5] $\times$ 36.0 14.4
PONI[19] $\times$ 31.8 12.1
ZSON[15] $\times$ 25.5 12.6 15.3 4.8
CoW[7] $\sqrt{}$ 7.4 3.7
ESC[32] $\sqrt{}$ 39.2 22.3 28.7 14.2
L3MVN[28] $\sqrt{}$ 50.4 23.1 36.3 15.7 34.9 14.5
OpenFMNav[9] $\sqrt{}$ 54.9 24.4 37.2 15.7
InstructNav[14] $\sqrt{}$ 58.0 20.9
VLFM[27]   $\sqrt{}\mkern-9.0mu{\smallsetminus}$ 52.5 30.4 63.6 32.5 36.4 17.5
VLFM*[27] $\sqrt{}$ 50.9 23.6 56.9 27.5 32.5 15.9
TriHelper[30] $\sqrt{}$ 56.5 25.3
SG-Nav[26] $\sqrt{}$ 54.0 24.9 49.6 25.5 40.2 16.0
ApexNav(Ours) $\sqrt{}$ 59.6 33.0 76.2 38.0 39.2 17.8
Figure 7: Failure Cause Statistics of ApexNav on the Three Datasets.
Implementation Details: In the Habitat setting, the agent selects actions from a discrete set $\mathcal{A}$: move_forward ($0.25\,\text{m}$), turn_left ($30^{\circ}$), turn_right ($30^{\circ}$), look_up ($30^{\circ}$), look_down ($30^{\circ}$), and stop. Each episode is limited to 500 steps. The agent operates within a perception range of $0.0\,\text{m},5.0\,\text{m}]$, with a success distance of $0.2\,\text{m}$. The onboard camera is mounted at a height of $0.88\,\text{m}$. For semantic segmentation, we use YOLOv7 [21], GroundingDINO [13], and Mobile-SAM [29]. The semantic score map is generated by BLIP-2 [10]. DeepSeek-V3 [11] is used to generate the similar object list, rooms where objects are likely to appear, and detection confidence thresholds. The parameters are set as follows: $r_{t}=1.10$, $\sigma_{t}=0.015$ (Sec. IV-B1), and $d_{\text{local}}=0.75$, $\omega_{t}=200$, $\omega_{p}=2000$, $\omega_{s}=1.0$, $\epsilon=10^{-2}$, $K=10$, $d_{\text{th}}=0.15$ (Sec. IV-D). All simulation experiments are run on a system with an NVIDIA GeForce RTX 4090 GPU and an Intel i9-13900K ×32 CPU.
Baselines: SemExp [5], PONI [19], and ZSON [15] rely on task-specific training, limiting their generalization in zero-shot settings. CoW [7] uses a nearest-frontier exploration strategy without semantic reasoning. ESC [32], L3MVN [28], TriHelper [30], and OpenFMNav [9] improve exploration by building semantic maps and using LLMs to select frontiers. VLFM [27] and InstructNav [14] use LLMs or VLMs to create value maps that guide agents. VLFM* replaces the pre-trained PointNav module with a shortest-path planner for fair comparison. SG-Nav [26] further enhances semantic reasoning by prompting LLMs with 3D scene graphs to select frontiers.
@@ V-B Comparison with State-of-the-art
V-C Failure Cause Analysis
Task outcomes are categorized as follows:
Except for the HM3Dv2 dataset, which mainly consists of single-floor tasks, the others show many ”Different Floor” failures ($13\%$). This is because ApexNav builds only a single-layer 2D map, so such cross-floor tasks are treated as failures by default. This limitation also leads to many ”No Frontier” failures. When the agent starts at a staircase between two floors, the vertical distance to the target is not large enough to be recognized as on a different floor. However, since the agent builds a 2D map centered at its starting position, it may find no frontiers to explore after initialization, resulting in task failure. This limitation is mainly an engineering constraint, and could be addressed by incorporating 3D maps in future work.
@@ V-C Failure Cause Analysis
V-D Ablation Study on Adaptive Exploration Strategy
V-D1 Ablation Study on Exploration Strategy
We conduct an ablation study on HM3Dv2 to evaluate how different exploration strategies affect performance. Four strategies were tested: (a) TSP-Semantic: selects high-scoring frontiers and solves TSP to plan the order; (b) Greedy-Semantic: chooses the frontier with the highest semantic score; (c) TSP-Geometry: solves TSP over all frontiers; (d) Greedy-Geometry: selects the nearest frontier. (e) Utility-based: selects the highest utility score (semantic score divided by path length [6]). We evaluate single-stage strategies (using one method throughout), two-stage strategies (switching between semantic and geometric modes based on thresholds $\sigma_{t}=0.015$ and $r_{t}=1.10$) and utility-based strategies. Results are shown in Table II.
The results show that two-stage strategies outperform single-stage ones, highlighting the advantage of adaptively switching between semantic and geometric exploration. Among the single-stage strategies, TSP-Semantic and Greedy-Geometry performed best, and their combination achieves the highest overall score: Greedy-Geometry is efficient when semantic cues are weak, while TSP-Semantic guides the agent effectively once strong semantic cues emerge. Meanwhile, our strategy outperforms utility-based methods by better prioritizing semantically informative regions, even when they are far away, unlike utility-based methods that often undervalue such areas due to longer path costs.
@@ V-D2 Ablation Study on Exploration Metrics
(a) w/o Similar Objects 56.95 30.40 75.00 37.63 28.70 15.43
(b) w/o Fusion 44.50 25.48 55.00 29.72 27.84 14.03
(c) w/o Fusion (High DCT) 55.20 31.24 68.20 34.99 37.81 17.16
(d) w/o Observation 50.48 29.17 70.20 37.44 35.58 16.09
(e) w Max-fusion 51.43 28.97 67.90 37.04 32.11 15.32
@@
V-E Ablation Study on Target-centric Semantic Fusion
To assess the impact of the target-centric semantic fusion method, we conduct many ablation experiments. The following control groups were designed: (a) w/o Similar Objects: Removes the auxiliary detection module for similar objects; (b, c) w/o Fusion: Removes the semantic fusion module and directly overwrites cluster confidence using detection results; (b) uses a detector confidence threshold (DCT) as in ApexNav, while (c) uses a higher DCT as in VLFM. A lower DCT makes the detector more prone to false positives. (d) w/o Observation: The confidence of the object cluster is not reduced by fusing low-confidence detection results when no objects are detected; (e) w Max-fusion: Replaces ApexNav’s fusion method with max-confidence fusion. The results are presented in Table IV.
Both the (b) and (c) groups underperform compared to Ours, emphasizing the importance of fusion. Although (c) reduces false detections with a higher DCT, it also misses some correct ones. ApexNav uses a lower DCT and refines more candidates through fusion, leading to better results and showing stronger robustness to false detections. Additionally, ApexNav outperforms the max-fusion approach (e), proving its effectiveness in preventing task failures caused by high-confidence mislabeling. The other two experiments emphasize the significance of fusing similar objects and reducing confidence for objects that were previously detected but are currently undetected. Overall, the target-centric semantic fusion method delivers the best performance, demonstrating its effectiveness.
V-F Ablation Study on Safe Waypoint Navigation
To validate the efficiency and safety of ApexNav’s navigation strategy, we conduct an ablation study by removing the safe waypoint navigation component and relying solely on the shortest path computed using the A* algorithm for navigation. SR, SPL, and DTG were chosen as evaluation metrics. The experimental results are summarized in Table V.
[…]
The Shortest Path 66.20 33.85 1.72
Safe Navigation (Ours) 76.20 38.03 1.31
TABLE V: Ablation Study on Safe Waypoint Navigation. This table compares the result with ”The Shortest Path” on the HM3Dv2 dataset.
The results reveal that relying solely on the shortest path for navigation leads to a notable decline in SR and SPL, accompanied by a significant increase in DTG. This indicates that the agent’s traversal paths deteriorate, often resulting in collisions that disrupt stable progress toward the target. These findings highlight the effectiveness of the safe waypoint navigation method in enhancing both efficiency and safety.
V-G Real-world Deployment
We conduct real-world experiments using the AgileX LIMO robot. A laptop equipped with an RTX 4060 Laptop GPU and an AMD R9 7945HX CPU served as the ROS master and communicated with the robot via a ROS-based master-slave architecture. We use MINCO [22] to generate collision-free spatiotemporal trajectories for waypoint navigation. ApexNav is evaluated in dormitory and laboratory environments across six trials, targeting toilet, couch, and chair in the former, and refrigerator, plant, and TV in the latter. Representative results and quantitative statistics are shown in Fig. 1 and Table VII. Module runtimes are summarized in Table VI. Most modules run within the desired interval, even on a consumer-grade laptop. Although detection slightly exceed this range, it has little effect on overall task performance. This further highlights the efficiency and deployability in practical scenarios.
Modules GPU Runtime (ms) Desired
RTX 4060 Laptop RTX 4090 Interval (ms)
Detection+Segmentation $[190,320]$ $[65,125]$ 250
BLIP2 $[160,190]$ $[50,55]$ 200
(a) Runtime of modules on different GPU devices.
Modules CPU Runtime (ms) Desired
AMD Ryzen9 7945hx Intel i9-13900K Interval (ms)
Semantic Object Map $[15,20]$ $[5,15]$ 250
Semantic Score Map $[1,2]$ $[0.5,1]$ 200
Grid Map $[30,50]$ $[20,35]$ 100
ESDF Map $[1,3]$ $[1,2]$ 100
Exploration Planning $[50,200]$ $[40,150]$ 500
(b) Runtime of modules on different CPU devices.
TABLE VI: Time consumption of modules on different devices. The Desired Interval indicates the runtime range required for stable real-time performance.
Trials SR (%) SPL TD (m) TT (s)
6 100 0.723 15.85 56
TABLE VII: Quantitative Statistics in Real-world Experiments Results are averaged over all trials. TD: Traveled Distance; $T$: Traveled Time.